OpenStudy (jazzyfa30):

Determine the slope of the line that passes through the points at (1, 5) and (-4, 0). Select one: a. -1 b. 0 c. 5/3 d. 1

OpenStudy (jazzyfa30):

@mathslover help

sam (.sam.):

Take the coordinates \((x_1,y_1)\) and \((x_2,y_2)\) Slope formula is \[Slope=\frac{y_2-y_1}{x_2-x_1}\] Substitute and find slope

OpenStudy (jazzyfa30):

\[\frac{ 1+4 }{5-0}=\frac{ 5 }{ 5 }=1\]

OpenStudy (phi):

yes, 1 - (-4) is 1+4= 5 5-0 is 5 5/5= 1

OpenStudy (phi):

But you should be careful. the top is the y-values: 5-0 and the x values are in the bottom 1 - (-4) in this case you get 5/5, but in general you would get the wrong answer if you put the x on top

OpenStudy (phi):

Try to remember "change in Y divided by change in X"

OpenStudy (jazzyfa30):

ok thanks @phi can you help me with a few more